OpenStudy (anonymous):

1. If I know the lengths of two sides of a right triangle, how do I find the third? 2. Could I find the two missing side lengths of a right triangle if I only know one side length and one angle measure (other than the 90 degree angle)? 3. Could I find the two missing angle measures if I know some of the side lengths of a right triangle? 4. What makes a triangle a “special” right triangle? How can special right triangles help me find side lengths?

OpenStudy (lgbasallote):

1. pythagorean theorem (it depends on what you're missing..pythagorean theorem is c^2 = a^2 + b^2 where a and b are the legs and c is the hypotenuse) 2. yes you can...you use sin/cos/tan 3. yes you can...you use sin/cos/tan 4. special right triangles are the 30-60-90 and the 45-45 90 in 30-60-90 triangles the legs are always in the form of this: hypotenuse = twice the length of the side opposite 30 side opposite 30 = 1/2 of the hypotenuse side opposite 60 = the length of the side opposite 60 x sqrt 3 in 45-45-90 the sides opposite 45 are equal and they are equal to sqrt 2 times the hypotenuse not sure about the last sentence...but i know that they are equal and there is a sqrt of 2 somewhere...just dunno if it's in the hypotenuse or the legs